Weight: 131 lbs

Female

Height: 5 ft 2 inches

Calculate, showing the math, your Body Mass Index (BMI) using the formula 'weight in kilograms divided by height in meters squared' (kg/m^2). Note: Weight in pounds divided by 2.2 = weight in kg and height in inches times .0254 = height in meters. The number is absolute and not a percentage.

Based on your BMI, are you underweight, healthy weight, overweight, or obese?
Answer #1

Weight = 131 lbs = 131 × 0.454 = 59.47 kg

Height = 5 feet 2 imches = 5.2 × 30.48 = 158.5kg

BMI = Weight ÷ (Height in meters)² = 59.47÷ (158.5)²

= 23.67

- Body mass index (BMI):

- A BMI of less than 18.5 means that a person is underweight.

- A BMIof between 18.5 and 24.9 is ideal.

- A BMI of between 25 and 29.9 is overweight.

- A BMI over 30 indicates obesity.

Base on this, it is a healthy weight.